OpenStudy (anonymous):

Simplify 2.5*(x+0.6)^2 = 3.5x^2

OpenStudy (anonymous):

\[2.5*(x+0.6)^2 = 3.5x^2\]

OpenStudy (anonymous):

do you know the laws of the indices

OpenStudy (anonymous):

Thus far I've simplified it to \[(x+0.6)^2=1.4x^2\] however, I'm not sure whether I should square root at this point. I do know the laws of the indices so I know I can't distribute the square over the parenthesis.

OpenStudy (anonymous):

expand the left hand side

OpenStudy (anonymous):

\[x^2+1.2x+.36=1.4x^2\] then solve a quadratic equation,